1/4(y+8)+6=10 whats y and how do you show your work
Answer by waynest(281)

!/4(y + 8) + 6 = 10
y/4 + (1/4 x 8) + 6 = 10
y/4 + 2 + 6 = 10
y/4 + 8 = 10
y/4 + 8 - 8 = 10 - 8
y/4 = 2
4(y/4) = 4(2)
y = 8
check:
1/4(8 + 8) + 6 = 10
1/4(16) + 6 = 10
4 + 6 = 10
10 = 10